### v1

# TReMu: Towards Neuro-Symbolic Temporal Reasoning for LLM-Agents with Memory in Multi-Session Dialogues

Abstract
Temporal reasoning in multi-session dialogues presents a significant challenge which has been under-studied in previous temporal reasoning benchmarks. To bridge this gap, we propose a new evaluation task for temporal reasoning in multi-session dialogues and introduce an approach to construct a new benchmark by augmenting dialogues from LoCoMo and creating multi-choice QAs. Furthermore, we present TReMu, a new framework aimed at enhancing the temporal reasoning capabilities of LLM-agents in this context. Specifically, the framework employs time-aware memorization through timeline summarization, generating retrievable memory by summarizing events in each dialogue session with their inferred dates. Additionally, we integrate neuro-symbolic temporal reasoning, where LLMs generate Python code to perform temporal calculations and select answers. Experimental evaluations on popular LLMs demonstrate that our benchmark is challenging, and the proposed framework significantly improves temporal reasoning performance compared to baseline methods, raising from 29.83 on GPT-4o via standard prompting to 77.67 via our approach and highlighting its effectiveness in addressing temporal reasoning in multi-session dialogues.
TReMu: Towards Neuro-Symbolic Temporal Reasoning for LLM-Agents with Memory in Multi-Session Dialogues
Yubin Ge1111thanks: Work done as an intern at AWS, Salvatore Romeo2222, Jason Cai2222, Raphael Shu2222, Monica Sunkara2222, Yassine Benajiba2222, Yi Zhang2222 $1$University of Illinois Urbana Champaign, USA $2$Amazon Web Services yubinge2@illinois.edu {romeosr, cjinglun, sunkaral, zhongzhu, sunkaral, yizhngn}@amazon.com
1 Introduction
In the context of multi-session dialogues, temporal reasoning is both critical and challenging for LLM-agents. As the number of dialogue sessions increases, storing and retrieving relevant information efficiently becomes more difficult Maharana et al. (2024). LLMs often face challenges in managing large, long-term dialogues, such as failing to retrieve specific temporal details from long history and dialogues exceed the input limit of LLMs. Additionally, research has shown that LLMs overlook important contextual information from long dialogue histories due to the accumulation of irrelevant historical data, referred to as "historical noise" Wang et al. (2023). These challenges underscore the need for enhanced temporal reasoning capabilities in LLM-agents for effective handling of multi-session dialogues.
Figure 1: Examples from LoCoMo showing the two temporal characteristics we focus on in this work.
However, most existing temporal reasoning benchmarks cannot be used directly for this study, because they are built on shorter texts, such as stories and Wikipedia articles, that contain clear temporal information Chen et al. (2021); Wang and Zhao (2024); Xiong et al. (2024). Even benchmarks designed for dialogues, like TimeDial Qin et al. (2021) and LoCoMo Maharana et al. (2024), do not explicitly account for the special temporal characteristics in multi-session dialogues, such as relative time and cross-session dependency. For instance, speakers often use relative time expressions instead of specific dates, requiring the model to infer exact event times. Moreover, it is common for speakers to recall past events from previous sessions, creating cross-session dependencies, where events from different sessions involve the same or related entities and reflect changes over time. This further requires LLMs to retain context effectively when reasoning about events across multiple sessions.
In this work, we present TReMu (Temporal Reasoning for LLM-Agents in Multi-Session Dialogues), a novel framework designed to enhance temporal reasoning in multi-session dialogues. Our framework introduces time-aware memorization, which uses timeline summarization to generate summaries for each dialogue session, identifying events and associating them with their inferred dates. These summaries, linked to specific times (either session times or inferred event dates), serve as retrievable memory. This effectively addresses events expressed in relative time by distinguishing when such an event occurred from when it was mentioned by the speaker.
During reasoning, we propose a neuro-symbolic temporal reasoning approach inspired by recent work that integrates LLMs with symbolic reasoning, translating questions into symbolic language before using a solver to find answers Pan et al. (2023); Olausson et al. (2023). Specifically, given a temporal question, we retrieve relevant memory and instruct the LLMs to generate Python code. This approach leverages the LLMs’ strong Python coding capabilities and existing Python libraries for temporal calculations. The generated code serves as an intermediate rationale. By executing the code line-by-line, the model follows step-by-step reasoning similar to CoT Wei et al. (2022), leading the model to select the correct answer.
Due to the absence of temporal reasoning evaluation benchmarks specific to multi-session dialogues, we propose a method to construct a new evaluation benchmark focusing on two key temporal characteristics: relative time and cross-session dependency. By augmenting dialogues from LoCoMo Maharana et al. (2024), we create multiple-choice temporal questions spanning three types of reasoning to evaluate the temporal reasoning capabilities of LLMs in this context.
We evaluate our framework based on three popular LLMs—GPT-4o, GPT-4o-mini, and GPT-3.5-Turbo—on our benchmark. The results show that our benchmark is challenging, revealing suboptimal performance for LLMs. In contrast, our framework demonstrates superior performance compared to baseline methods, such as CoT, highlighting the effectiveness of our approach in improving temporal reasoning in multi-session dialogues.
Our contributions are as follows:
  • We propose a new framework for temporal reasoning in multi-session dialogues, integrating time-aware memorization and neuro-symbolic temporal reasoning.
  • We propose a method to construct a temporal reasoning evaluation benchmark for multi-session dialogues by augmenting an existing dataset, explicitly covering the temporal characteristics of relative time and cross-session dependency.
  • Through extensive experiments, we empirically show that temporal reasoning in multi-session dialogues poses significant challenges for LLMs, even with strategies like CoT. However, our framework significantly improves LLMs’ temporal reasoning in this context.
Dialogue Dataset Avg. Turns Per Conv. Avg. Sessions Per Conv. Avg. Tokens Per Conv. Time Interval Collection
MSC Xu et al. (2022) 53.3 4 1,225.9 few days Crowdsourcing
Conversation Chronicles Jang et al. (2023) 58.5 5 1,054.7 few hours - years LLM-generated
LoCoMo Maharana et al. (2024) (Ours) 304.9 19.3 9,209.2 few months LLM-gen. + crowdsourcing
Table 1: Statistics of the chosen multi-session dialogue dataset, LoCoMo, compared to others.
2 Benchmark Construction
In this section, we introduce the construction pipeline to build our temporal QA benchmark for evaluating LLM-agents’ temporal reasoning in multi-session dialogues. As mentioned earlier, we mainly focus on the two temporal characteristics in multi-session dialogues: relative time and cross-session dependency.
Question Type Count ##\## of Options ##\## of Events Event Type
Temporal Anchoring 264 5 1 relative time
Temporal Precedence 102 3 2 cross session dependency (+ relative time)
Temporal Interval 234 5 2 cross session dependency (+ relative time)
Total 600
  Total - Unanswerable 112
LoCoMo Maharana et al. (2024) 321
Table 2: Dataset statistics and details of the constructed benchmark.##\## of Options refers to the number of options for each temporal question. ##\## of Events refers to the number of selected events to create each temporal question. Event Type specifies the type of temporal events chosen for question creation, where (+ relative time) indicates that relative time was an additional consideration in event selection. Note that our benchmark not only contains more temporal QAs than LoCoMo, but also include unanswerable questions.
2.1 Benchmark Design
We propose augmenting an existing multi-turn dialogue dataset to create a benchmark for evaluating LLM-agents’ temporal reasoning in multi-session dialogues. After a thorough examination, we selected LoCoMo Maharana et al. (2024), which comprises dialogues averaging 600 turns and 16,000 tokens across up to 32 sessions. In comparison to existing multi-session dialogue datasets, LoCoMo features the longest dialogues and the most sessions, thus presenting a greater challenge. The detailed statistics of LoCoMo and its comparison with other relevant datasets are shown in Table 1.
As mentioned earlier, our benchmark focuses on two key temporal characteristics in multi-session dialogues: relative time and cross-session dependency. To achieve this, we follow previous benchmarks Chen et al. (2021); Xiong et al. (2024); Wang and Zhao (2024) by creating temporal QA pairs based on temporal events in the dialogues. Specifically, we design each temporal QA based on either a single event or a pair of events:
$\bullet$ Single Event: We select events expressed with relative time and develop a temporal reasoning type called Temporal Anchoring, which asks for the exact time of the event.
$\bullet$ Two Events: We choose pairs of relevant events from different sessions that exhibit cross-session dependency. We also consider relative time as an extra factor to increase the complexity of the questions. Two temporal reasoning types are applied: Temporal Precedence, which asks which event occurred first, and Temporal Interval, which asks for the duration between the two events.
2.2 Construction Pipeline
To construct our benchmark, we follow the design of our benchmark and utilize a systematic step-by-step approach with GPT-4o. The prompt for each step is shown in Appendix Sec.A.
Step 1: Temporal Event Extraction We begin by prompting GPT-4o to extract all temporal events from each dialogue session. In addition, we instruct GPT-4o to annotate the relative time expressions for these events, facilitating the selection process for creating temporal QAs.
Step 2: Temporal Event Linking Next, we link the extracted events containing cross-session dependency within the dialogue. We prompt GPT-4o with the extracted events and instruct it to group those related to the same or relevant entities across different sessions, particularly those reflecting changes in attributes over time. For example, the event “Debra Ryan told her mentor about her business idea” from an early session is linked to “Debra Ryan started her own business” from a later session.
Step 3: Temporal QA Creation Once the temporal events are processed, we prompt GPT-4o to select those events that meet the criteria for different temporal reasoning types and generate multiple-choice temporal QAs. Additionally, we create unanswerable questions, as in prior QA benchmarks Rajpurkar et al. (2018), to comprehensively assess models’ temporal reasoning capabilities.
Step 4: Quality Control We observe various noises in generated QAs, such as incorrect inferences about exact times or the selection of events with ambiguous time expressions (e.g., "the other day"). To ensure the benchmark’s quality, we follow recent temporal reasoning benchmarks for LLMs, such as TGQA Xiong et al. (2024), to perform quality control. We manually review each question to verify that it aligns with our design specifications and that the answers are correctly grounded in the dialogue. We also revise well-constructed questions with incorrect answers and remove any unreasonable ones. The final temporal QA benchmark covers time intervals from days to months and its statistics and details are presented in Table 2. Particularly, our final benchmark not only contains more temporal QAs than LoCoMo, but also include unanswerable questions, which are not covered in LoCoMo. We also include examples of QAs for different temporal reasoning types in Appendix Sec. B.
3 Methodology
3.1 Preliminary: Memory-Augmented LLM-Agents
To address the limit of LLMs struggling in retaining information from long input text, recent studies have demonstrated that equipping LLM agents with memory can efficiently support long-turn conversations Lu et al. (2023); Packer et al. (2023); Zhong et al. (2024). Therefore, we base our study on memory-augmented LLM-agents.
The general pipeline of memory-augmented LLM-agents comprises three stages: memorization, retrieval, and response. In the memorization stage, the model summarizes each dialogue session and stores these summaries as memory. During the retrieval stage, the model retrieves the most relevant memory for the current dialogue session. This retrieved memory is then concatenated with the ongoing dialogue to generate the next response.
For temporal reasoning in multi-session dialogues, we can apply this pipeline by summarizing each dialogue session, retrieving relevant memory for each temporal question, and using this information to select the answer. Specifically, we build our framework based on MemoChat Lu et al. (2023), which effectively realizes this three-stage process through prompting and has demonstrated effectiveness in handling long-range dialogues.
3.2 TReMu
Building on the memory-augmented LLM-agent pipeline, we introduce our framework called TReMu as shown in Algorithm 1. The framework consists of two key components: time-aware memorization and neuro-symbolic temporal reasoning.
Algorithm 1 TReMu
  Initialize:
  Time-aware Memorization Model LLMmem
  Memory Retrieval Model LLMretrieval
  Neuro-symbolic Reasoning Model LLMcode
  Python Executor $\mathcal{P}$
  Memorization pool $\mathcal{M}\leftarrow\emptyset$
  {Time-aware Memorization}
  for each dialogue session $d_{i}$ in dialogue $\mathcal{D}$ do
     $m_{i}\leftarrow\text{LLM}_{\text{mem}}(d_{i})$
     $\mathcal{M}\leftarrow\mathcal{M}\cup m_{i}$
  end for
  {Neuro-symbolic Temporal Reasoning}
  for each temporal question $q$ do
     $m_{retrieved}\leftarrow\text{LLM}_{\text{retrieval}}(q,\mathcal{M})$
     $c\leftarrow\text{LLM}_{code}(q,m_{retrieved})$
     $o\leftarrow\mathcal{P}(c)$
     final answer $a\leftarrow\text{LLM}(q,o)$
  end for
3.2.1 Time-aware Memorization
As LLMs may struggle with capturing temporal information in dialogues, particularly relative time, we propose a time-aware memorization strategy using timeline summarization Steen and Markert (2019); Rajaby Faghihi et al. (2022); Sojitra et al. (2024). Specifically, for each dialogue session tagged with its timestamp, we prompt the LLM to generate its summary while also summarizing mentioned events occurring on different inferred dates. The prompt is shown in Appendix Sec.C. This allows us to obtain a series of summaries for a single dialogue session, with each summary is linked to a specific date. By distinguishing between when an event occurred and when it was mentioned, this time-aware memorization could improve temporal reasoning and enable more precise event tracking across the dialogue sessions.
3.2.2 Neuro-symbolic Temporal Reasoning
Allen’s interval algebra Allen (1983) provides a foundation for solving various types of temporal reasoning by identifying event timestamps and performing temporal comparisons or calculations. For example, determining the difference between the timestamps of two events allows us to resolve temporal interval problems. Building on the interval algebras, previous works have developed various neuro-symbolic models to integrate symbolic reasoning with neural networks Garcez and Lamb (2003); Zhou et al. (2021).
Additionally, researchers have explored neuro-symbolic reasoning for LLMs. By providing LLMs with clear instructions about the grammar of the symbolic language and offering a few demonstrations as in-context examples, LLMs can accurately translate problems into different symbolic language, such as first-order logic Han et al. (2022). Once the problem is translated, symbolic solvers can be employed to solve it, and it has been demonstrated to outperform methods like CoT for multi-hop reasoning tasks Pan et al. (2023).
Inspired by the above motivations, we propose leveraging LLMs to translate temporal reasoning questions into Python code, which can then be executed to derive answers. The detailed prompts are shown in Appendix Sec.D. The underlying reasons stem from the strong ability of state-of-the-art LLMs in generating Python code in multiple code generation benchmarks and by the availability of Python libraries, such as datetime and dateutil, which support temporal calculations. Particularly, dateutil provides a function relativedelta supporting relative time calculation, for example we can infer next Friday using TODAY + relativedelta(weekday=FR). Meanwhile, we provide some implemented functions that the model can directly call, such as "weekRange(t)" returns the start date and the end date of the week that t lies in. As LLMs have familiarity with Python, we provide demonstration via in-context learning to guide it in generating Python code for temporal reasoning, based on the given question and retrieved memory. We then execute the output Python code, and the LLM is prompted again to select the correct answer by using the code.
Particularly, we believe our reasoning approach offers an alternative form of CoT. While the original CoT Wei et al. (2022) performs step-by-step reasoning in natural language, our neuro-symbolic approach conducts temporal reasoning by executing generated code line-by-line in a programming language. This neuro-symbolic method retains the core concept of CoT’s step-by-step reasoning while leveraging the precision and additional support provided by Python code and packages, enhancing the accuracy of the reasoning process.
4 Experiments
4.1 Baselines and Metrics
We build our framework using various backbone LLMs: GPT-4o [^1] Specifically, gpt-4o-2024-05-13., GPT-4o-mini [^2] Specifically, gpt-4o-mini-2024-07-18., and GPT-3.5-Turbo [^3] Specifically, gpt-3.5-turbo-0125.. Particularly, for GPT-3.5-Turbo, many of LoCoMo dialogues are longer than its input length, we then follow LoCoMo Maharana et al. (2024) which earlier dialogues are omitted.
Particularly, in our setting of multi-session dialogues, where conversations may exceed the input limits of LLMs, we also consider the memory mechanism as a critical component of baselines in order to feed complete dialogue information. Therefore, following recent research on temporal reasoning with LLMs Wang and Zhao (2024); Xiong et al. (2024), we compare our approach with the following baseline methods, including direct prompting and LLM-agents with memory:
Methods Accuracy Unanswerable Questions
TA TP TI Overall Precision Recall F1
SP 18.18 58.82 30.34 29.83 46.88 13.39 20.84
CoT 67.80 74.51 49.15 61.67 42.61 43.75 43.18
MemoChat 35.23 43.14 25.21 32.67 24.30 77.68 37.02
Memochat + CoT 51.14 49.02 26.50 41.67 24.80 81.25 38.00
Timeline + CoT 83.33 78.41 58.55 71.50 48.51 58.04 52.84
TReMu 84.47 81.37 68.38 77.67 55.48 76.79 64.42
Table 3: Experimental results of various methods based on GPT-4o. We use TA to represent Temporal Anchoring, TP for Temporal Precedence and TI for Temporal Interval.
Methods Accuracy Unanswerable Questions
TA TP TI Overall Precision Recall F1
SP 20.08 50.00 29.91 29.00 40.00 26.79 32.08
CoT 46.59 62.75 37.18 45.67 33.96 48.21 39.86
MemoChat 21.21 39.22 23.50 25.17 21.11 74.11 32.88
Memochat + CoT 24.62 45.10 24.36 28.00 21.11 75.00 32.94
Timeline + CoT 55.68 59.80 38.46 49.67 30.73 59.82 40.60
TReMu 64.02 46.08 38.89 51.17 29.21 92.86 44.44
Table 4: Experimental results of various methods based on GPT-4o-mini.
Methods Accuracy Unanswerable Questions
TA TP TI Overall Precision Recall F1
SP 21.59 31.37 23.08 23.83 22.91 46.43 30.68
CoT 23.86 38.24 22.65 25.83 20.97 50.00 29.56
MemoChat 17.42 45.10 23.50 24.50 21.93 66.96 33.04
Memochat + CoT 20.45 53.92 26.50 28.50 21.79 50.00 30.36
Timeline + CoT 32.58 44.12 22.65 30.67 22.57 51.79 31.44
TReMu 42.42 37.25 22.22 33.67 23.33 75 35.60
Table 5: Experimental results of various methods based on GPT-3.5-Turbo.
  • Standard Prompting (SP): The entire dialogue is provided along with each temporal question, with additional instructions for selecting the correct answer.
  • Chain-of-Thought (CoT) Wei et al. (2022): Similar to SP, but with additional instructions for LLMs to solve questions step-by-step.
  • MemoChat Lu et al. (2023): Given that multi-turn dialogues can exceed the model’s input length, and since our approach builds on memory-augmented LLM-agents, MemoChat serves as a baseline where we modify the response stage to answer temporal questions.
  • MemoChat + CoT: This baseline applies CoT in the response stage to answer temporal questions step-by-step using the retrieved memory.
  • Timeline + CoT: Based on the framework of memory-augmented LLM-agents, we modify the original memorization with our proposed timeline summarization and combine it with CoT as a baseline.
For evaluation metrics, we primarily use accuracy to assess the overall performance of temporal reasoning. In addition, for unanswerable questions, we calculate precision, recall, and the F1 score to specifically measure performance on this subset of questions. Specifically, precision is computed as the accuracy of questions the model predicts as "unanswerable," while recall is determined by the accuracy of questions where the ground truth answer is "unanswerable."
4.2 Experimental Results
The results are shown in Tables 3, 4, and 5 for GPT-4o, GPT-4o-mini, and GPT-3.5-Turbo, respectively. On the recent benchmark TRAM Wang and Zhao (2024), existing LLMs have shown strong performance via direct prompting. For instance, on TRAM, GPT-4 achieves an accuracy of 82 with CoT, and GPT-3.5 reaches 71.40 using CoT. In contrast, our benchmark poses a greater challenge, as GPT-4o achieves only 61.67 with CoT and 29.83 with SP, while GPT-3.5’s performance is significantly lower, with 25.83 using CoT and 23.83 with SP. These gaps are likely due to the challenging context of multi-session dialogues and the temporal characteristics in multi-session dialogues, which are not explicitly covered by previous benchmarks.
Besides, our proposed framework outperforms all baseline methods across all three LLMs in terms of both accuracy and F1 score, with a notable increase in accuracy from 29.83 with SP to 77.67 with our framework using GPT-4o. This demonstrates the effectiveness of our approach in enhancing temporal reasoning for multi-session dialogues. A comparison between MemoChat + CoT and Timeline + CoT reveals that the form of memory matters. Time-aware memorization brings higher accuracy due to timeline summarization, which requires models to infer temporal information during memorization, particularly regarding relative time. Notably, incorporating the memory mechanism performs worse than CoT on GPT-4o and GPT-4o-mini. This may be because these models’ input lengths are sufficient to handle LoCoMo dialogues, allowing them to identify relevant temporal information without memory augmentation. In contrast, for GPT-3.5, which has a shorter input limit, adding the memory mechanism generally improves performance by enabling the model to retrieve information from complete dialogues rather than truncated ones.
Additionally, including CoT generally improves the performance, consistent with previous findings Wang and Zhao (2024); Xiong et al. (2024). However, changing from CoT to symbolic reasoning, as implemented in our framework, yields further improvements. This can be attributed to the models’ proficiency in generating Python code without losing the advantage of step-by-step reasoning. When models accurately translate questions into Python, the execution of the code avoids errors, offering an advantage over purely neural reasoning. These results align with recent research that combines LLMs with symbolic reasoners for various reasoning tasks Olausson et al. (2023); Pan et al. (2023).
Figure 2: The percentage of execution failures.
Figure 3: The first case study comparing CoT and our proposed framework, where CoT results in the wrong answer "E" but our approach selects the correct option "C". We highlight the key information in colors.
Figure 4: The second case study comparing CoT and our proposed framework, where CoT results in the wrong answer "E" but our approach selects the correct option "C". We highlight the key information in colors.
4.3 Execution Failure Study
We observe that LLMs occasionally generate Python code with syntax errors, such as typos or mismatched brackets. To delve into this, we measure the percentage of generated code that fails to execute and, during inference, we regenerate the code when such errors occur. The results, shown in Figure 2, indicate that the percentages of execution failure are generally low across all three LLMs, demonstrating the reliability of our Python-based symbolic reasoning approach. As expected, GPT-4o exhibits the lowest rate of execution failure, while GPT-3.5-Turbo has the highest, corresponding to the overall performance differences we demonstrate above in temporal reasoning among these models. This likely reflects the inherent performance gap between the LLMs.
4.4 Case Study
In this section, we demonstrate how the two key components of our framework—time-aware memorization and neuro-symbolic temporal reasoning—work in real cases. We compare the outputs of CoT and our framework based on GPT-4o.
In Figure 3, with CoT, even though GPT-4o successfully identifies that the key temporal information is "Sharon’s survival course started on 12 March 2020," but it gets confused with "week-long course" and infers the end date of the course, incorrectly selecting "Unanswerable." In contrast, with our framework’s time-aware memorization, the model retrieves the event from memory along with its properly inferred time. During the reasoning stage, the model utilizes this memory to distinguish between when the speaker, Sharon, mentioned the event (03/16/2020) and when the event occurred (03/12/2020). Then the model defines the corresponding variable in the generated code, i.e., t__\__start__\__course, to precisely capture the time.
Figure 4 illustrates another mistake made via CoT. The model correctly infers that the "last week" corresponds to the session time of 16 March 2020 but fails to match the week range with the correct answer—the week of 03/09/2020 is the week of 03/11/2020 but the model does not realize this. As for our framework, the model leverages the Python dateutil package’s relativedelta function, alongside our custom weekRange function, to accurately infer the last week’s range. This neuro-symbolic reasoning not only facilitates the model to reason step-by-step but also enhances it by incorporating external temporal functions to support more accurate temporal reasoning.
5 Related Work
Temporal Reasoning for LLMs.
Recent advancements in large language models (LLMs) have brought significant improvements in reasoning capabilities Huang and Chang (2023), leading to growing interest in temporal reasoning Chu et al. (2024); Qiu et al. (2024). Existing approaches primarily address this challenge through time-aware language modeling. For example, Kanashiro Pereira; Tan et al. propose fine-tuning strategies to enhance temporal reasoning, while Zhou et al.; Yang et al. introduce auxiliary objectives to incorporate external temporal knowledge. However, studies such as Chu et al.; Qiu et al. show that state-of-the-art LLMs still exhibit suboptimal performance in temporal reasoning with prompting techniques. Our framework differs from these works by utilizing memory-augmented LLM agents, enhancing memorization through timeline summarization, and integrating neuro-symbolic reasoning as an intermediate step for answering temporal questions.
Multi-session Dialogues.
Several studies have focused on multi-session dialogues and developed benchmarks to address this task. Xu et al. introduced MSC, the first multi-session dataset incorporating time intervals between sessions. Similarly, Bae et al. proposed a dynamic memory management method to maintain up-to-date user information and introduced a Korean multi-session dialogue dataset. Jang et al. created the CONVERSATION CHRONICLES dataset, designed for long-term conversations that integrate time intervals and detailed speaker relationships. More recently, Maharana et al. introduced LoCoMo, a dataset featuring long-term dialogues spanning up to 35 sessions. While our work is situated within this context, it specifically targets temporal reasoning, addressing the temporal characteristics of relative time and cross-session dependency, which have not been explicitly explored in prior research.
6 Conclusion
In this paper, we address the critical challenge of temporal reasoning in multi-session dialogues for LLM-agents. We present a method to construct a temporal reasoning evaluation benchmark by augmenting dialogues from LoCoMo, specifically targeting the temporal characteristics of relative time and cross-session dependency. Furthermore, we introduce a novel framework which combines time-aware memorization through timeline summarization with neuro-symbolic temporal reasoning by translating temporal questions into executable Python code. Through extensive evaluations, we demonstrate that our benchmark presents significant challenges, and our framework substantially outperforms baseline methods in improving temporal reasoning for multi-session dialogues.
7 Limitations
Our work has several limitations. First, our constructed temporal reasoning benchmark consists of multiple-choice questions, while real multi-session dialogues involve users interacting with LLM agents in a generative manner. A benchmark that incorporates generative temporal reasoning questions could more accurately reflect the agents’ reasoning abilities in practical scenarios. We plan to extend our current benchmark to support generative question-answering in future work.
Additionally, we observe that the LoCoMo dialogues contain some temporal ambiguities that can affect reasoning over related events. For instance, if a speaker mentions on a Tuesday that an event occurred "last Monday," it is unclear whether the event happened the previous day or on the Monday of the prior week. However, in real applications, LLM agents could potentially resolve such ambiguities by asking follow-up clarifying questions.
Appendix A Prompts for Benchmark Construction
We use GPT-4o to construct a temporal reasoning benchmark for multi-session dialogues. The first step is the temporal event extraction using the prompt shown in Figure 5. Then the prompt for the second step, temporal event linking, is shown in Figure 6. With the grouped temporal events, we use the prompt in Figure 7 to create temporal QAs.
Figure 5: Prompt for Temporal Event Extraction.
Figure 6: Prompt for Temporal Event Linking.
Figure 7: Prompt for Temporal QA Creation.
Appendix B Examples of Temporal QAs in Constructed Benchmark
We show examples of final temporal QAs for different temporal reasoning types in Figure 8, 9 and 10. In each example, we highlight the ground truth answer as green and show the corresponding selected temporal events for constructing the question below the question.
Figure 8: An example temporal QA for Temporal Anchoring
Figure 9: An example temporal QA for Temporal Precedence
Figure 10: An example temporal QA for Temporal Interval
Appendix C Prompt for Time-aware Memorization
We show the designed prompt for time-aware memorization in Figure 11.
Figure 11: Prompt for Time-aware Memorization.
Appendix D Prompt for Neuro-symbolic Temporal Reasoning
We show the designed prompts for neuro-symbolic temporal reasoning here. Specifically, we first perform memory retrieval with the prompt in Figure 12. Then we prompt to generate Python code via in-context learning as in Figure 13. With the generated code and its execution result, we finally prompt the LLM to select the answer, and the prompt is shown in Figure 14.
Figure 12: Prompt for memory retrieval.
Figure 13: Prompt for generating Python code for temporal reasoning.
Figure 14: Prompt for temporal question answering.

### v2

@@
 Additionally, we integrate neuro-symbolic temporal reasoning, where LLMs generate Python code to perform temporal calculations and select answers.
-Experimental evaluations on popular LLMs demonstrate that our benchmark is challenging, and the proposed framework significantly improves temporal reasoning performance compared to baseline methods, raising from 29.83 on GPT-4o via standard prompting to 77.67 via our approach and highlighting its effectiveness in addressing temporal reasoning in multi-session dialogues.
+Experimental evaluations on popular LLMs demonstrate that our benchmark is challenging, and the proposed framework significantly improves temporal reasoning performance compared to baseline methods, raising from 29.83 on GPT-4o via standard prompting to 77.67 via our approach and highlighting its effectiveness in addressing temporal reasoning in multi-session dialogues.N
 TReMu: Towards Neuro-Symbolic Temporal Reasoning for LLM-Agents with Memory in Multi-Session Dialogues
-Yubin Ge1111thanks: Work done as an intern at AWS, Salvatore Romeo2222, Jason Cai2222, Raphael Shu2222, Monica Sunkara2222, Yassine Benajiba2222, Yi Zhang2222 $1$University of Illinois Urbana Champaign, USA $2$Amazon Web Services yubinge2@illinois.edu {romeosr, cjinglun, sunkaral, zhongzhu, sunkaral, yizhngn}@amazon.com
+Yubin Ge1,21,2thanks: Work done as an intern at AWS, Salvatore Romeo22, Jason Cai22, Raphael Shu22, Yassine Benajiba22thanks: Work done while at AWS, now at Oracle, Monica Sunkara22, Yi Zhang22 11University of Illinois Urbana Champaign, USA 22Amazon Web Services {yubinge, romeosr, cjinglun, zhongzhu, sunkaral, yizhngn}@amazon.com
 1 Introduction
+Temporal reasoning, a fundamental cognitive ability, plays a crucial role in human perception Hoerl and McCormack (2019).
+It encompasses not only basic concepts such as ordering and duration but also more complex functions like task planning and causal relation discovery Vila (1994).
+Besides, large language models (LLMs) have brought significant advancements to various tasks.
+These models typically undergo pre-training on extensive text corpora, endowing them with unprecedented accuracy in predicting the next token given some context Ge et al.
+(2023a).
+Despite the remarkable advancements in LLMs and their emerging reasoning capabilities Huang and Chang (2023), previous studies have shown that their performance in temporal reasoning remains suboptimal Chu et al.
+(2024); Qiu et al.
+(2024).
 In the context of multi-session dialogues, temporal reasoning is both critical and challenging for LLM-agents.
-As the number of dialogue sessions increases, storing and retrieving relevant information efficiently becomes more difficult Maharana et al.
-(2024).
-LLMs often face challenges in managing large, long-term dialogues, such as failing to retrieve specific temporal details from long history and dialogues exceed the input limit of LLMs.
+As dialogue sessions proceed, storing and retrieving relevant information efficiently becomes more difficult Maharana et al.
+(2024), such as failing to retrieve specific temporal details from long history and dialogues exceed the input limit of LLMs.
 Additionally, research has shown that LLMs overlook important contextual information from long dialogue histories due to the accumulation of irrelevant historical data, referred to as "historical noise" Wang et al.
@@
 Figure 1: Examples from LoCoMo showing the two temporal characteristics we focus on in this work.
-However, most existing temporal reasoning benchmarks cannot be used directly for this study, because they are built on shorter texts, such as stories and Wikipedia articles, that contain clear temporal information Chen et al.
+However, most existing temporal reasoning benchmarks cannot be used directly for this study, because they are usually built on shorter texts, such as stories and Wikipedia articles, that contain clear temporal information Chen et al.
 (2021); Wang and Zhao (2024); Xiong et al.
@@
 (2021) and LoCoMo Maharana et al.
-(2024), do not explicitly account for the special temporal characteristics in multi-session dialogues, such as relative time and cross-session dependency.
+(2024), do not explicitly consider the special temporal characteristics in multi-session dialogues, such as relative time and cross-session dependency.
 For instance, speakers often use relative time expressions instead of specific dates, requiring the model to infer exact event times.
@@
 As mentioned earlier, we mainly focus on the two temporal characteristics in multi-session dialogues: relative time and cross-session dependency.
-Question Type Count ##\## of Options ##\## of Events Event Type
+Question Type Count #\# of Options #\# of Events Event Type
 Temporal Anchoring 264 5 1 relative time
@@
 LoCoMo Maharana et al. (2024) 321
-Table 2: Dataset statistics and details of the constructed benchmark.##\## of Options refers to the number of options for each temporal question.
-##\## of Events refers to the number of selected events to create each temporal question.
+Table 2: Dataset statistics and details of the constructed benchmark.#\# of Options refers to the number of options for each temporal question.
+#\# of Events refers to the number of selected events to create each temporal question.
 Event Type specifies the type of temporal events chosen for question creation, where (+ relative time) indicates that relative time was an additional consideration in event selection.
@@
 2.1 Benchmark Design
-We propose augmenting an existing multi-turn dialogue dataset to create a benchmark for evaluating LLM-agents’ temporal reasoning in multi-session dialogues.
+We propose augmenting an existing dataset to create a benchmark for evaluating LLM-agents’ temporal reasoning in multi-session dialogues.
 After a thorough examination, we selected LoCoMo Maharana et al.
 (2024), which comprises dialogues averaging 600 turns and 16,000 tokens across up to 32 sessions.
-In comparison to existing multi-session dialogue datasets, LoCoMo features the longest dialogues and the most sessions, thus presenting a greater challenge.
-The detailed statistics of LoCoMo and its comparison with other relevant datasets are shown in Table 1.
+In comparison to existing multi-session dialogue datasets, LoCoMo features the longest dialogues and the most sessions (as shown in Table 1), thus presenting a greater challenge.
 As mentioned earlier, our benchmark focuses on two key temporal characteristics in multi-session dialogues: relative time and cross-session dependency.
@@
 2.2 Construction Pipeline
-To construct our benchmark, we follow the design of our benchmark and utilize a systematic step-by-step approach with GPT-4o. The prompt for each step is shown in Appendix Sec.A.
+To construct our benchmark, we follow the design of our benchmark and utilize a systematic step-by-step approach with GPT-4o. The prompt for each step is shown in Appendix $\S$ A.
 Step 1: Temporal Event Extraction We begin by prompting GPT-4o to extract all temporal events from each dialogue session.
@@
 (2018), to comprehensively assess models’ temporal reasoning capabilities.
-Step 4: Quality Control We observe various noises in generated QAs, such as incorrect inferences about exact times or the selection of events with ambiguous time expressions (e.g., "the other day").
+Step 4: Quality Control We observe various noises in generated QAs, such as incorrect inferences about exact times.
 To ensure the benchmark’s quality, we follow recent temporal reasoning benchmarks for LLMs, such as TGQA Xiong et al.
@@
 Particularly, our final benchmark not only contains more temporal QAs than LoCoMo, but also include unanswerable questions, which are not covered in LoCoMo.
-We also include examples of QAs for different temporal reasoning types in Appendix Sec.
+We also include examples of QAs for different temporal reasoning types in Appendix $\S$.
 B.
@@
 3.1 Preliminary: Memory-Augmented LLM-Agents
-To address the limit of LLMs struggling in retaining information from long input text, recent studies have demonstrated that equipping LLM agents with memory can efficiently support long-turn conversations Lu et al.
+To address the limit of LLMs struggling in retaining information from long input text, recent studies turn to equip LLM agents with memory to support long-turn conversations Lu et al.
 (2023); Packer et al.
@@
 This retrieved memory is then concatenated with the ongoing dialogue to generate the next response.
-For temporal reasoning in multi-session dialogues, we can apply this pipeline by summarizing each dialogue session, retrieving relevant memory for each temporal question, and using this information to select the answer.
 Specifically, we build our framework based on MemoChat Lu et al.
-(2023), which effectively realizes this three-stage process through prompting and has demonstrated effectiveness in handling long-range dialogues.
+(2023), which realizes this three-stage process through prompting and has demonstrated effectiveness in handling long-range dialogues.
 3.2 TReMu
@@
  Neuro-symbolic Reasoning Model LLMcode
-  Python Executor $\mathcal{P}$
+ Symbolic Solver $\mathcal{P}$
  Memorization pool $\mathcal{M}\leftarrow\emptyset$
@@
   $m_{i}\leftarrow\text{LLM}_{\text{mem}}(d_{i})$
-     $\mathcal{M}\leftarrow\mathcal{M}\cup m_{i}$
+  $\mathcal{M}\leftarrow f_{org}(\mathcal{M},m_{i})$
 end for
@@
 3.2.1 Time-aware Memorization
-As LLMs may struggle with capturing temporal information in dialogues, particularly relative time, we propose a time-aware memorization strategy using timeline summarization Steen and Markert (2019); Rajaby Faghihi et al.
+Our time-aware memorization builds on timeline summarization Steen and Markert (2019); Rajaby Faghihi et al.
 (2022); Sojitra et al.
-(2024).
-Specifically, for each dialogue session tagged with its timestamp, we prompt the LLM to generate its summary while also summarizing mentioned events occurring on different inferred dates.
-The prompt is shown in Appendix Sec.C.
-This allows us to obtain a series of summaries for a single dialogue session, with each summary is linked to a specific date.
-By distinguishing between when an event occurred and when it was mentioned, this time-aware memorization could improve temporal reasoning and enable more precise event tracking across the dialogue sessions.
+(2024) and it consists of two steps: Temporal Memory Writing and Memory Organization.
+During Temporal Memory Writing (prompt in Appendix $\S$.C), we instruct LLM agents to generate memory pieces while also extracting and associating mentioned events with inferred dates.
+Unlike prior approaches that summarize entire sessions holistically, our method produces fine-grained memory pieces linked to specific inferred time markers.
+As shown in Tables 3 and 4, our memorization outputs memory pieces corresponding to events with inferred time steps that facilitates to mitigate temporal ambiguity.
+For example, the highlighted texts show that Michelle cooked a meal and later referenced cooking it at different dates.
+This enables finer temporal granularity, effectively distinguishing events based on inferred time intervals.
+Topic Summary
+Catching Up Daniel and Michelle catch up on new events in their lives including new jobs, hobbies, and activities.
+Hobbies and Daily Rituals Daniel and Michelle discuss their hobbies and rituals like running, ballet, playing guitar, meditation, and cooking.
+Cooking and Celebrations Both talk about their cooking experiences and celebrate Daniel’s promotion.
+Books and Recommendations Michelle and Daniel discuss books they’ve read and recommend some to each other.
+Personal Items with Sentimental Value Michelle and Daniel talk about cameras and a vintage motorcycle with sentimental value.
+Table 3: Output memory from MemoChat based on one dialogue session in LoCoMo.
+Time Summary
+01/28/2020 Daniel and Michelle share updates on their lives… including Michelle starting her Masters in Psychology, Daniel starting a new job where he learns to code and problem-solve, and Michelle’s hobby of ballet, meditation, and journaling… Michelle mentions she made an Italian meal last Saturday and Daniel made salsa for a taco night.
+Daniel also shared receiving a promotion …
+01/27/2020 Daniel received a promotion and celebrated with a dinner at his favorite spot.
+01/25/2020 Michelle cooked a delicious Italian meal for her friends, including pasta, garlic bread, and tiramisu.
+01/24/2020 Daniel made a huge batch of salsa and hosted a taco night with friends.
+01/20/2020 Michelle started her Masters in Psychology.
+Table 4: Output memory from Time-aware Memorization based on the same dialogue session in LoCoMo.
+Then, we perform Memory Organization on the output memory pieces to maintain long-term memory.
+We structure memory in a timeline format, grouping events that occur simultaneously and indexing them based on inferred timesteps.
+This approach enhances the distinction between an event’s occurrence and its mention, reducing temporal ambiguity and improving time-based retrieval.
+These enhancements mark a significant difference from traditional memorization approaches, supporting efficiency in temporal reasoning.
 3.2.2 Neuro-symbolic Temporal Reasoning
-Allen’s interval algebra Allen (1983) provides a foundation for solving various types of temporal reasoning by identifying event timestamps and performing temporal comparisons or calculations.
-For example, determining the difference between the timestamps of two events allows us to resolve temporal interval problems.
-Building on the interval algebras, previous works have developed various neuro-symbolic models to integrate symbolic reasoning with neural networks Garcez and Lamb (2003); Zhou et al.
-(2021).
-Additionally, researchers have explored neuro-symbolic reasoning for LLMs.
-By providing LLMs with clear instructions about the grammar of the symbolic language and offering a few demonstrations as in-context examples, LLMs can accurately translate problems into different symbolic language, such as first-order logic Han et al.
-(2022).
-Once the problem is translated, symbolic solvers can be employed to solve it, and it has been demonstrated to outperform methods like CoT for multi-hop reasoning tasks Pan et al.
-(2023).
-Inspired by the above motivations, we propose leveraging LLMs to translate temporal reasoning questions into Python code, which can then be executed to derive answers.
-The detailed prompts are shown in Appendix Sec.D.
-The underlying reasons stem from the strong ability of state-of-the-art LLMs in generating Python code in multiple code generation benchmarks and by the availability of Python libraries, such as datetime and dateutil, which support temporal calculations.
+Inspired by the recent progress in neuro-symbolic reasoning for LLMs Han et al.
+(2022); Pan et al.
+(2023), we propose leveraging LLMs to translate temporal reasoning questions into Python code as intermediate rationales, which is executed as the reasoning process to derive answers (prompts shown in Appendix $\S$.D).
+We tried different symbolic languages and finally chose Python because SOTA LLMs are better at generating Python code and there exist Python libraries that support temporal calculations, like datetime and dateutil.
 Particularly, dateutil provides a function relativedelta supporting relative time calculation, for example we can infer next Friday using TODAY + relativedelta(weekday=FR).
-Meanwhile, we provide some implemented functions that the model can directly call, such as "weekRange(t)" returns the start date and the end date of the week that t lies in.
-As LLMs have familiarity with Python, we provide demonstration via in-context learning to guide it in generating Python code for temporal reasoning, based on the given question and retrieved memory.
-We then execute the output Python code, and the LLM is prompted again to select the correct answer by using the code.
-Particularly, we believe our reasoning approach offers an alternative form of CoT.
+Meanwhile, we provide implemented functions to be directly called, such as "weekRange(t)" returns the start date and the end date of the week that t lies in.
+Different from other works in temporal reasoning based on code executionLi et al.
+(2023), we enable our LLM agents with function-calling abilities, ensuring correctness and expanding the range of temporal reasoning tasks beyond simple precedence relations.
+We provide demonstration via in-context learning to generate Python code with function calling, given then question and retrieved memory.
+After the generated code is executed, the output and code serve as intermediate rationales, and the LLM is prompted again to give the answer.
+Particularly, our reasoning approach offers an alternative form of CoT.
 While the original CoT Wei et al.
 (2022) performs step-by-step reasoning in natural language, our neuro-symbolic approach conducts temporal reasoning by executing generated code line-by-line in a programming language.
-This neuro-symbolic method retains the core concept of CoT’s step-by-step reasoning while leveraging the precision and additional support provided by Python code and packages, enhancing the accuracy of the reasoning process.
-4 Experiments
-4.1 Baselines and Metrics
-We build our framework using various backbone LLMs: GPT-4o [^1] Specifically, gpt-4o-2024-05-13., GPT-4o-mini [^2] Specifically, gpt-4o-mini-2024-07-18., and GPT-3.5-Turbo [^3] Specifically, gpt-3.5-turbo-0125..
-Particularly, for GPT-3.5-Turbo, many of LoCoMo dialogues are longer than its input length, we then follow LoCoMo Maharana et al.
-(2024) which earlier dialogues are omitted.
-Particularly, in our setting of multi-session dialogues, where conversations may exceed the input limits of LLMs, we also consider the memory mechanism as a critical component of baselines in order to feed complete dialogue information.
-Therefore, following recent research on temporal reasoning with LLMs Wang and Zhao (2024); Xiong et al.
-(2024), we compare our approach with the following baseline methods, including direct prompting and LLM-agents with memory:
+This neuro-symbolic method retains the core concept of CoT’s step-by-step reasoning while leveraging the precision and additional support provided by Python code and packages.
+However, prior works Li et al.
+(2023) rely solely on solver outputs without providing intermediate justifications.
 Methods Accuracy Unanswerable Questions
@@
 Table 7: Experimental results of various methods based on GPT-3.5-Turbo.
+4 Experiments
+4.1 Experimental Setup
+Models.
+We build our framework using various black-box LLMs: GPT-4o [^1] Specifically, gpt-4o-2024-05-13., GPT-4o-mini [^2] Specifically, gpt-4o-mini-2024-07-18., and GPT-3.5-Turbo [^3] Specifically, gpt-3.5-turbo-0125..
+Particularly, for GPT-3.5-Turbo, many of LoCoMo dialogues are longer than its input length, we then follow LoCoMo Maharana et al.
+(2024) which earlier dialogues are omitted.
+Particularly, we have also tried different open-source LLMs but most of them cannot handle the long dialogue inputs from LoCoMo, for example only about 10% of dialogues can be fed into Llama-3-70B.
+And even for those shorter dialogues that can be fed into Llama-3-70B, we notice that the model gets lost and fails to follow instructions, even failing to generate in the desired format.
+Therefore, we leave the adaptation to LLMs as future work.
+Baselines.
+Since in our setting of multi-session dialogues where the conversations exceed the input limits of LLMs, we consider the memory mechanism as a critical component of baselines in order to feed complete dialogue information.
+Therefore, we include the following baselines for comparison:
   • Standard Prompting (SP): The entire dialogue is provided along with each temporal question, with additional instructions for selecting the correct answer.
@@
 (2023): Given that multi-turn dialogues can exceed the model’s input length, and since our approach builds on memory-augmented LLM-agents, MemoChat serves as a baseline where we modify the response stage to answer temporal questions.
+To better understand the effectiveness of each component in our framework, we evaluate the following variants as baselines for the ablation study:
   • MemoChat + CoT: This baseline applies CoT in the response stage to answer temporal questions step-by-step using the retrieved memory.
   • Timeline + CoT: Based on the framework of memory-augmented LLM-agents, we modify the original memorization with our proposed timeline summarization and combine it with CoT as a baseline.
-For evaluation metrics, we primarily use accuracy to assess the overall performance of temporal reasoning.
+Comparing MemoChat + CoT and Timeline + CoT allows us to assess the impact of replacing the standard memory mechanism in LLM agents with our time-aware memorization.
+Additionally, comparing Timeline + CoT with TReMu highlights the effect of replacing CoT with our neuro-symbolic reasoning approach.
+Evaluation Metrics.
+We primarily use accuracy to assess the overall performance of temporal reasoning.
 In addition, for unanswerable questions, we calculate precision, recall, and the F1 score to specifically measure performance on this subset of questions.
@@
 4.2 Experimental Results
-The results are shown in Tables 3, 4, and 5 for GPT-4o, GPT-4o-mini, and GPT-3.5-Turbo, respectively.
-On the recent benchmark TRAM Wang and Zhao (2024), existing LLMs have shown strong performance via direct prompting.
-For instance, on TRAM, GPT-4 achieves an accuracy of 82 with CoT, and GPT-3.5 reaches 71.40 using CoT.
-In contrast, our benchmark poses a greater challenge, as GPT-4o achieves only 61.67 with CoT and 29.83 with SP, while GPT-3.5’s performance is significantly lower, with 25.83 using CoT and 23.83 with SP.
-These gaps are likely due to the challenging context of multi-session dialogues and the temporal characteristics in multi-session dialogues, which are not explicitly covered by previous benchmarks.
-Besides, our proposed framework outperforms all baseline methods across all three LLMs in terms of both accuracy and F1 score, with a notable increase in accuracy from 29.83 with SP to 77.67 with our framework using GPT-4o.
+The results are shown in Tables 5, 6, and 7 for GPT-4o, GPT-4o-mini, and GPT-3.5-Turbo, respectively.
+On the recent TRAM benchmark Wang and Zhao (2024), existing LLMs demonstrate strong performance with direct prompting.
+For instance, GPT-4 achieves an accuracy of 82 using CoT, while GPT-3.5 attains 71.40.
+In contrast, our benchmark is significantly more challenging.
+GPT-4o achieves only 61.67 with CoT and 29.83 with SP, whereas GPT-3.5 performs even worse, scoring 25.83 with CoT and 23.83 with SP.
+These performance gaps likely stem from the complexity of multi-session dialogues and their temporal dependencies, which are not explicitly addressed in previous benchmarks.
+Our framework outperforms all baseline methods across all three LLMs in terms of both accuracy and F1 scores, with a notable increase in accuracy from 29.83 with SP to 77.67 with our framework using GPT-4o.
 This demonstrates the effectiveness of our approach in enhancing temporal reasoning for multi-session dialogues.
-A comparison between MemoChat + CoT and Timeline + CoT reveals that the form of memory matters.
-Time-aware memorization brings higher accuracy due to timeline summarization, which requires models to infer temporal information during memorization, particularly regarding relative time.
-Notably, incorporating the memory mechanism performs worse than CoT on GPT-4o and GPT-4o-mini.
-This may be because these models’ input lengths are sufficient to handle LoCoMo dialogues, allowing them to identify relevant temporal information without memory augmentation.
-In contrast, for GPT-3.5, which has a shorter input limit, adding the memory mechanism generally improves performance by enabling the model to retrieve information from complete dialogues rather than truncated ones.
-Additionally, including CoT generally improves the performance, consistent with previous findings Wang and Zhao (2024); Xiong et al.
+However, incorporating a memory mechanism performs worse than CoT for GPT-4o and GPT-4o-mini.
+This may be because these models have sufficient input lengths to process LoCoMo dialogues, enabling them to identify relevant temporal information without additional memory augmentation.
+In contrast, for GPT-3.5, which has a shorter input limit, the memory mechanism generally improves performance by allowing the model to retrieve information from memory rather than truncated dialogues.
+Furthermore, we find that incorporating CoT generally improves performance, aligning with previous findings Wang and Zhao (2024); Xiong et al.
 (2024).
-However, changing from CoT to symbolic reasoning, as implemented in our framework, yields further improvements.
-This can be attributed to the models’ proficiency in generating Python code without losing the advantage of step-by-step reasoning.
-When models accurately translate questions into Python, the execution of the code avoids errors, offering an advantage over purely neural reasoning.
-These results align with recent research that combines LLMs with symbolic reasoners for various reasoning tasks Olausson et al.
-(2023); Pan et al.
-(2023).
+In particular, CoT encourages models to search for relevant information within dialogues.
+However, due to the dialogues’ length, the models sometimes generate responses like "I cannot find the mention of …," which hinders their temporal reasoning capabilities.
+This further underscores the necessity of a memory mechanism to support long dialogue settings.
 Figure 2: The percentage of execution failures.
@@
 We highlight the key information in colors.
+4.3 Ablation Study
+From Tables 5, 6, and 7, the comparison between MemoChat + CoT and Timeline + CoT highlights the importance of memory representation.
+Time-aware memorization improves accuracy by instructing models to infer temporal information—particularly relative time—during memorization and mitigate temporal ambiguity.
+Furthermore, replacing CoT with symbolic reasoning, as seen in the comparison between Timeline + CoT and TReMu, leads to additional performance gains.
+This improvement stems from the models’ ability to generate Python code while retaining the benefits of step-by-step reasoning.
+This aligns with recent research integrating LLMs with symbolic reasoners for various reasoning tasks Olausson et al.
+(2023); Pan et al.
+(2023).
 4.4 Execution Failure Study
-We observe that LLMs occasionally generate Python code with syntax errors, such as typos or mismatched brackets.
-To delve into this, we measure the percentage of generated code that fails to execute and, during inference, we regenerate the code when such errors occur.
+We also measure the percentage of generated code that fails to execute and, during inference, we regenerate the code when such errors occur.
 The results, shown in Figure 2, indicate that the percentages of execution failure are generally low across all three LLMs, demonstrating the reliability of our Python-based symbolic reasoning approach.
@@
 During the reasoning stage, the model utilizes this memory to distinguish between when the speaker, Sharon, mentioned the event (03/16/2020) and when the event occurred (03/12/2020).
-Then the model defines the corresponding variable in the generated code, i.e., t__\__start__\__course, to precisely capture the time.
+Then the model defines the corresponding variable in the generated code, i.e., t_\_start_\_course, to precisely capture the time.
 Figure 4 illustrates another mistake made via CoT.
@@
 Multi-session Dialogues.
-Several studies have focused on multi-session dialogues and developed benchmarks to address this task.
+Except from the massive research efforts on common dialogue generation Ge et al.
+(2023b), several studies have developed multi-session dialogues benchmarks.
 Xu et al.
@@
 More recently, Maharana et al.
-introduced LoCoMo, a dataset featuring long-term dialogues spanning up to 35 sessions.
+introduced LoCoMo, a dataset featuring long-term and multi-modal dialogues.
 While our work is situated within this context, it specifically targets temporal reasoning, addressing the temporal characteristics of relative time and cross-session dependency, which have not been explicitly explored in prior research.
@@
 In this paper, we address the critical challenge of temporal reasoning in multi-session dialogues for LLM-agents.
-We present a method to construct a temporal reasoning evaluation benchmark by augmenting dialogues from LoCoMo, specifically targeting the temporal characteristics of relative time and cross-session dependency.
+We present a method to construct a temporal reasoning evaluation benchmark by augmenting dialogues from LoCoMo and covering the temporal characteristics of relative time and cross-session dependency.
 Furthermore, we introduce a novel framework which combines time-aware memorization through timeline summarization with neuro-symbolic temporal reasoning by translating temporal questions into executable Python code.
@@
 7 Limitations
-Our work has several limitations.
-First, our constructed temporal reasoning benchmark consists of multiple-choice questions, while real multi-session dialogues involve users interacting with LLM agents in a generative manner.
-A benchmark that incorporates generative temporal reasoning questions could more accurately reflect the agents’ reasoning abilities in practical scenarios.
-We plan to extend our current benchmark to support generative question-answering in future work.
-Additionally, we observe that the LoCoMo dialogues contain some temporal ambiguities that can affect reasoning over related events.
-For instance, if a speaker mentions on a Tuesday that an event occurred "last Monday," it is unclear whether the event happened the previous day or on the Monday of the prior week.
-However, in real applications, LLM agents could potentially resolve such ambiguities by asking follow-up clarifying questions.
+Our work has several limitations:
+Assessment in QA settings.
+Our evaluation follows the standard practice of recent temporal reasoning benchmarks like Zhou et al.
+(2019); Wang and Zhao (2024), using a multiple-choice format for reliable assessment.
+A more ideal setting would be evaluating in generative dialogue settings.
+However, generative QA evaluations pose significant challenges due to variations in temporal expressions (e.g., "January 1st" vs.
+"01/01"), making exact match and F1 token score unreliable.
+Thus, we prioritize benchmark reliability and accuracy over a more ambitious generative QA setting.
+We leave the extension of the current benchmark to generative dialogue settings as future work
+Open-sourced Models.
+Our current experiments mainly focus on close-sourced models.
+However, as we have pointed out in our Experiment section $\S$ 4.1, we found most open-source LLMs cannot handle the long dialogue inputs from LoCoMo, for example only about 10% of dialogues can be fed into Llama-3-70B.
+And even for those shorter dialogues that can be fed into Llama-3-70B, we notice that the model gets lost and fails to follow instructions, even failing to generate in the desired format.
+We therefore consider the adaptation of our framework to open-sourced models as future work.
+Note that we will release our code and dataset for reproducibility.
 Appendix A Prompts for Benchmark Construction